Jim can fill a pool carrying bucks of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together?
Answer by Yoan(9)

First thing you need to do is get them to equal the same thing. Which is one job. they all added will equal to 1.
Jim can fill a pool carrying bucks of water in 30 minutes so think of it as 1/30 X (X is a unit of time). Sue can do the same job in 45 minutes which is 1/45 X. Tony can do the same job in 1 ½ hours lets change that to minutes which is 90 minutes so it would be 1/90 X. How quickly can all three fill the pool together? here it is 1/30 X + 1/45 X + 1/90 X = 1
we take the GCD which is 90 so it is
3/90 X + 2/90 X + 1/90 X = 1
6/90 X = 1
X = 90/6
X = 15
they do it in 15 minuts